# Dual-scale Enhanced and Cross-generative Consistency Learning for Semi-supervised Polyp Segmentation

### v1

@@ IV-B Comparison with State-of-the-arts
We compare the proposed DEC-Seg with ten state-of-the-art semi-supervised segmentation methods, i.e., MT [15], DAN [34], UA-MT [10], URPC [16], Duo-SegNet [32], CLCC [17], SLC-Net [18], MC-Net+ [37], CDMA [38] and BCP [44]. For a fair comparison, we change the backbones of all compared methods to “Res2Net”, and then train all compared methods without any data augmentation. Moreover, we adopt a basic U-Net [19] model with Res2Net as the backbone of the encoder, denoted as “Baseline”.
IV-B1 Quantitative Results
Quantitative comparisons are reported in Tab. I and Tab. II. To validate the learning ability of the two seen training datasets (CVC-ClinicDB and Kvasir), we can be seen from Tab. I that while most methods learn well, our method performs better than other compared methods, both on $10\%$ and $30\%$ of labeled data. In addition, due to the additional unlabeled data, our method improves mDice and mIoU by $8.2\%$ and $12.0\%$ compared to baseline using only $10\%$ labeled data, and by $4.9\%$ and $6.3\%$ compared to baseline using only $30\%$ labeled data on CVC-ClinicDB, which indicate that our DEC-Seg can fully leverage unlabeled data to improve the polyp segmentation performance.
@@ IV-B1 Quantitative Results
Additionally, we report the comparison results on three unseen datasets (CVC-300, ETIS, and CVC-ColonDB) from Tab. II to verify the generalization ability of our DEC-Seg. Because the polyp in the CVC-300 dataset is similar to the training set and relatively simple, the metrics are also similar to CVC-ClinicDB and Kvisir datasets. Compared with SLC-Net, which performs relatively stable on this dataset, our method achieves $3.0\%$ and $3.5\%$ improvements in terms of mDice and mIoU with $30\%$ labeled data. The CVC-ColonDB and ETIS datasets are more difficult, as polyps are with large variations in size. As reported in Tab. II, comparing the BCP method on the CVC-ColonDB dataset using $10\%$ labeled data, our method achieves $1.2\%$ and $2.1\%$ improvements on mDice and mIoU, respectively. Besides, our method achieves $5.3\%$ and $6.2\%$ improvements over the SOTA semi-supervised segmentation method (BCP) in terms of mDice and mIoU with the $30\%$ labeled data. Similarly, on the ETIS dataset, our performance has a more significant superiority over other methods. The main reason is that our exploration and utilization of scale information make the model more robust to scale variation, which is able to identify small polyps that cannot be identified by other methods, and can also refine the edges of large polyps. Overall, the results indicate that our DEC-Seg has a better generalization ability and effectively utilizes unlabeled data to boost polyp segmentation.
@@
Figure 6: Qualitative results of our model and other ten compared semi-supervised segmentation methods.
IV-B2 Qualitative Results
Fig. 6 shows some of the visualization results of the test examples under $30\%$ labeled data. It can be seen that our method can accurately locate and segment polyps under different challenging factors. For example, in the first two rows of Fig. 6, the polyps have very small sizes. It can be clearly seen that Duo-SegNet recognizes the polyp incorrectly, while UAMT, CLCC, MC-Net+, and CDMA methods even fail to locate polyps. Other methods only locate some fragments of polyps, while our method accurately and completely segments polyps. In the $3^{rd}$ and $4^{th}$ rows, the polyps are visually embedded in their surrounding mucosa, thus it is very difficult to accurately locate and segment these polyps. From the results, our method performs better than other comparison methods to accurately segment polyps. In the $5^{th}$ and $6^{th}$ rows, the polyps have relatively large sizes, making it is challenging to complete locate the polyps. In this case, some methods (e.g., URPC, Duo-SegNet, MC-Net+, and BCP) produce several regions with over-segmented fragments, while our method can obtain promising segmentation results and produce fine details of the boundary. This is mainly because our method makes use of multi-scale information and cross-generative consistency to learn more powerful feature representations, which are helpful to boost the segmentation performance. In addition, we also show some segmentation results containing multiple polyps (see the $7^{th}$ and $8^{th}$ rows). Compared with other methods, the segmentation of multiple polyps by our method is more complete and accurate. It can be also observed that our method can effectively locate and segment polyps under different challenging factors, such as scale variation, homogeneous regions, non-sharp boundaries, and multiple polyps.
TABLE III: Ablative results on the Kvasir and CVC-ColonDB datasets with $30\%$ labeled data.
@@ IV-B2 Qualitative Results
(No.1) Baseline 0.843 0.782 0.829 0.877 0.551 0.489 0.543 0.732
(No.2) Baseline + SC 0.859 0.801 0.850 0.886 0.600 0.538 0.593 0.757
(No.3) Baseline + SC + SPC 0.865 0.810 0.856 0.891 0.611 0.548 0.602 0.765
(No.4) Baseline + SC + SPC + CC 0.875 0.815 0.864 0.892 0.635 0.565 0.627 0.773
(No.5) Baseline + SC + SPC + CC + CFA 0.876 0.815 0.867 0.892 0.654 0.580 0.649 0.778
(No.6) Baseline + SC + SPC + CC + CFA + DCF (ours) 0.882 0.822 0.871 0.899 0.691 0.616 0.678 0.798
IV-C Ablation Study
@@
To verify the effectiveness of each key component in DEC-Seg, we conduct ablation studies with $30\%$ labeled data on Kvasir and CVC-ColonDB datasets. The ablative results are shown in Tab. III, where “Baseline” denotes the base segmentation framework only using the labeled data.
@@
Effectiveness of SC. To investigate the importance of scale-enhanced consistency (SC), we add the DC loss to encourage the consistency of predictions from the same inputs with different scales. From Tab. III, we observe that No.2 (Baseline + SC) outperforms No.1 and obtains an $8.9\%$ improvement in mean Dice. This result indicates that scale-enhanced consistency is very helpful to improve polyp segmentation performance and is robust to scale variation.
Effectiveness of SPC and CC. We further study the contributions of scale-aware perturbation consistency (SPC) and cross-generative consistency (CC). As shown in Tab. III, No.3 performs better than No.2, indicating the effectiveness of scale-aware perturbation consistency. Besides, it can be seen that No.4 improves the No.3 performance on the Kvasir dataset, as the mean Dice is improved from $0.865$ to $0.875$. Therefore, these improvements indicate that introducing cross-generative consistency loss can help accurately segment polyp tissues in the learning of details and textures.
Effectiveness of CFA: We then investigate the importance of the proposed CFA module. To achieve this, we add the CFA on the basis of No.4, by aggregating the features of adjacent layers in the encoder before feeding into the decoder. As reported in Tab. III, it can be clearly seen that No.5 has a significant improvement over No.4 on the CVC-ColonDB dataset with a large scale variation, which proves that CFA can effectively capture scale information and improve segmentation performance.
@@
Effectiveness of DCF. As shown in Tab. III, No.6 (using the proposed DCF module) outperforms No.5 on two datasets. This indicates that fusion of multi-scale features can further improve the segmentation performance. In addition, to further validate the effectiveness of the dual-scale complement fusion strategy in the proposed DCF module, we construct a “Basic” strategy, which conducts a concatenation operation followed by two convolution layers to integrate the two features from different scales. The comparison results are shown in Tab. IV. From Tab. IV, it can be observed that our DCF performs better than the “Basic” strategy, indicating the effectiveness of the designed dual-scale complementary fusion module. Moreover, we visualize the segmentation results by using three different decoders, i.e., the scale-specific and fused decoders, and the comparison results are shown in Fig. 7. It can be observed that our method can not accurately locate the boundaries of the polyps when only using the original scale features or the downsampled scale features. However, we integrate the features from the two scales and then propagate them into the fused decoder, which can produce more accurate segmentation maps (as shown in Fig. 7 (c)). This further validates that the fusion of the features from different scales is helpful for improving the segmentation performance.
@@
TABLE IV: Ablative results of our DCF module and a basic fusion strategy on the Kvasir and CVC-ColonDB datasets. The labeled ratio is set to $30\%$.
@@
Figure 7: Visualization results of the three decoders. (a) predictions of the decoder (i.e., $D_{s1}$, $D_{t1}$) using the original scale features; (b) predictions of the decoder (i.e., $D_{s2}$, $D_{t2}$) using the downsampled scale features; (c) predictions of the fused decoders (i.e., $D_{s}^{f}$, $D_{t}^{f}$) using the multi-scale integrated features. The red and blue lines denote the ground truth and predictions, respectively.
@@ V Conclusion
In this paper, we have presented a novel semi-supervised learning framework (DEC-Seg) for polyp segmentation with colonoscopy images. The proposed cross-level feature aggregation module integrates the adjacent features from different resolutions, to enhance the features’ representation ability. Then, a scale-enhanced consistency is proposed to handle scale variation and learn more scale-aware features. Meanwhile, we design the fused decoders and a dual-scale complementary fusion module to aggregate the features from the scale-specific decoders. Moreover, multiple consistency strategies, i.e., scale-aware perturbation consistency and cross-generative consistency, are presented to enhance the learning process and fully leverage unlabeled data to boost the segmentation performance. Further, we propose the dual-scale complementary fusion module to integrate the features from two scale-specific decoders, which are passed through the fused decoder to produce the final segmentation maps. Experimental results on the five benchmark datasets show that our DEC-Seg is superior to state-of-the-art semi-supervised segmentation methods.

### v2

@@ IV-C Comparison with State-of-the-arts
We compare the proposed DEC-Seg with ten state-of-the-art semi-supervised segmentation methods, i.e., MT [10], DAN [32], UA-MT [11], URPC [14], CLCC [15], SLC-Net [16], MC-Net+ [36], CDMA [37], SCP-Net [50] and MCF [51]. For a fair comparison, we change the backbones of all compared methods to “Res2Net”, and then train all compared methods without any data augmentation. All the compared methods undergo $10,000$ iterations, similar to our approach, utilizing the same data and image dimensions. In the case of CLCC, we maintain its original setting of $320\times 320$ due to the specificity of the method. Moreover, the optimizer, learning rate, and descent strategy all remain consistent with their original configurations across the compared methods.
IV-C1 Experiments on Polyp Segmentation
@@
Quantitative comparisons are reported in Tab. I and Tab. II. To validate the learning ability of the two seen training datasets (CVC-ClinicDB and Kvasir), it can be seen from Tab. I that while most methods learn well, our method performs better than other compared methods, both on $10\%$ and $30\%$ of labeled data, which indicate that our DEC-Seg can fully leverage unlabeled data to improve the polyp segmentation performance. Additionally, we report the comparison results on two unseen datasets (ETIS, and CVC-ColonDB) from Tab. II to verify the generalization ability of our DEC-Seg. The CVC-ColonDB and ETIS datasets are difficult, as polyps are with large variations in size. As reported in Tab. II, comparing the SCP-Net method on the CVC-ColonDB dataset using $30\%$ labeled data, our method achieves $3.9\%$ and $4.1\%$ improvements on mDice and mIoU, respectively. Similarly, on the ETIS dataset, our performance has a more significant superiority over other methods. The primary reason is that our exploration and utilization of scale information enhance the model’s robustness to scale variations. This capability allows it to small polyps, while other methods cannot identify them accurately. Overall, the results indicate that our DEC-Seg has a better generalization ability and effectively utilizes unlabeled data to boost polyp segmentation.
@@
Figure 6: Qualitative results of our model and other ten compared semi-supervised segmentation methods on polyp segmentation datasets.
Fig. 6 presents some of the visualization results of the test examples under $30\%$ labeled data. It can be seen that our method can accurately locate and segment polyps under different challenging factors. For example, in the first row of Fig. 6, the polyps have very small sizes. It can be seen that DAN, URPC, CLCC, and SLC-Net methods fail to locate polyps. Other methods produce over-segmented fragments and confuse the edges of polyps, while our method accurately and completely segments polyps. In the $3^{rd}$ row, the polyps are visually embedded in their surrounding mucosa, thus it is very difficult to accurately locate and segment these polyps. From the results, our method performs better than other comparison methods to accurately segment polyps. In the $4^{th}$ row, the polyps have relatively large sizes, making it challenging to complete locate the polyps. In this case, some methods (e.g., UA-MT, URPC, CLCC, CDMA, and MCF) only locate some fragments of polyps, while our method can obtain promising segmentation results and produce fine details of the boundary. This is mainly because our method makes use of multi-scale information and cross-generative consistency to learn more powerful feature representations, which help boost the segmentation performance. In addition, we also show some segmentation results containing multiple polyps (see the $5^{th}$ row). Compared with other methods, the segmentation of multiple polyps by our method is more complete and accurate. It can be also observed that our method can effectively locate and segment polyps under different challenging factors, such as scale variation, homogeneous regions, non-sharp boundaries, and multiple polyps.
[…]
Figure 7: Qualitative results of our model and other ten compared semi-supervised segmentation methods on brain tumor segmentation dataset.
[…]
Figure 8: Qualitative results of our model and other ten compared semi-supervised segmentation methods on skin lesion segmentation dataset.
IV-C2 Experiments on Brain Tumor Segmentation
As shown in Tab. III, our DEC-Seg achieves significant improvement in five metrics, and compared to the second best, our method achieves $5.2\%$ and $5.5\%$ improvements in terms of mDice and mIou with $10\%$ labeled data, and the corresponding improvements are $2.8\%$ and $4.8\%$ with $30\%$ labeled data. Moreover, it can be seen in Fig. 7 that our method can segment the glioma completely and coherently. In the $2^{rd}$ row, the segmentation maps of other methods are discrete, while our method produces more coherent segmented fragments. In addition, in the $4^{th}$ row, some methods (e.g., URPC, CLCC, and SLC-Net) do not find glioma regions, while our method accurately locates and segments glioma regions.
IV-C3 Experiments on Skin Lesion Segmentation
Tab. IV shows the performance on skin lesion segmentation dataset with $10\%$ and $30\%$ labeled ratios. DEC-Seg surpasses all state-of-the-arts. Our method achieves $1.3\%$ and $3.0\%$ improvements over MCF in terms of mDice and mIou with $10\%$ labeled data, and the corresponding improvements are $1.6\%$ and $3.0\%$ with $30\%$ labeled data. Besides, Fig. 8 shows the visual segmentation results in the skin lesion segmentation dataset. Compared with other methods, our DEC-Seg achieves better and more accurate segmentation in blurred boundaries.
TABLE V: Ablative results on the Kvasir and CVC-ColonDB datasets with $30\%$ labeled data.
@@ IV-C3 Experiments on Skin Lesion Segmentation
(No.1) Baseline 0.850 0.778 0.829 0.873 0.612 0.539 0.600 0.760
(No.2) Baseline + SC 0.873 0.812 0.865 0.894 0.637 0.572 0.632 0.773
(No.3) Baseline + SC + DCF 0.882 0.818 0.874 0.892 0.675 0.599 0.658 0.790
(No.4) Baseline + SC + DCF + CC 0.888 0.827 0.882 0.900 0.702 0.626 0.694 0.807
(No.5) Baseline + SC + DCF + CFA 0.888 0.824 0.878 0.898 0.698 0.621 0.688 0.804
(No.6) Baseline + SC + DCF + CC + CFA (ours) 0.893 0.830 0.886 0.902 0.721 0.640 0.709 0.814
IV-D Ablation Study
@@
To verify the effectiveness of each key component in DEC-Seg, we conduct ablation studies with $30\%$ labeled data on Kvasir and CVC-ColonDB datasets. The ablative results are shown in Tab. V, where “Baseline” denotes the semi-supervised framework with only scale-aware perturbation consistency.
@@
Effectiveness of SC. To investigate the importance of scale-enhanced consistency (SC), we add the SC loss to encourage the consistency of predictions from the same inputs with different scales. From Tab. V, we observe that No.2 (Baseline + SC) outperforms No.1 and obtains a $4.1\%$ improvement in mean Dice. This result indicates that scale-enhanced consistency is very helpful in improving polyp segmentation performance and is robust to scale variation.
@@
Effectiveness of DCF. As shown in Tab. V, No.3 (using the proposed DCF module) outperforms No.2 on two datasets. This indicates that the fusion of multi-scale features can further improve the segmentation performance. In addition, to further validate the effectiveness of the dual-scale complement fusion strategy in the proposed DCF module, we construct a “Basic” strategy, which conducts a concatenation operation followed by two convolution layers to integrate the two features from different scales. The comparison results are shown in Tab. VI. From Tab. VI, it can be observed that our DCF performs better than the “Basic” strategy, indicating the effectiveness of the designed dual-scale complementary fusion module. Moreover, we visualize the segmentation results by using three different decoders, i.e., the scale-specific and scale-fused decoders, and the comparison results are shown in Fig. 9. It can be observed that our method can not accurately locate the boundaries of the polyps when only using the original scale features or the downsampled scale features. However, we integrate the features from the two scales and then propagate them into the scale-fused decoder, which can produce more accurate segmentation maps (as shown in Fig. 9 (c)). This further confirms that integrating features from different scales enhances segmentation performance.
Effectiveness of CC. We further study the contributions of cross-generative consistency (CC). As shown in Tab. V, it can be seen that No.4 improves the No.3 performance on the CVC-ColonDB dataset, as the mean Dice is improved from $0.675$ to $0.702$. Therefore, these improvements indicate that introducing cross-generative consistency loss can help accurately segment polyp tissues in the learning of details and textures.
Effectiveness of CFA. We then examine the significance of the proposed CFA module. To do this, we integrate the CFA into configurations No.3 and No.4 by aggregating the features of adjacent layers in the encoder before passing them to the decoder. As shown in Tab. V, No.5 outperforms No.3, and No.6 shows a substantial improvement over No.4 on the CVC-ColonDB dataset, which has significant scale variation. This demonstrates that the CFA module effectively captures scale information and enhances segmentation performance.
@@
TABLE VI: Ablative results of our DCF module and a basic fusion strategy on the Kvasir and CVC-ColonDB datasets. The labeled ratio is set to $30\%$.
@@
Figure 9: Visualization results of the three decoders. (a) predictions of the decoder (i.e., $D_{1}$) using the original scale features; (b) predictions of the decoder (i.e., $D_{2}$) using the downsampled scale features; (c) predictions of the scale-fused decoders (i.e., $D_{f}$) using the multi-scale integrated features. The red and blue lines denote the ground truth and predictions, respectively.
[…]
Figure 10: Some failure cases generated by our DEC-Seg.
IV-E Limitation and Future Work
The qualitative and quantitative evaluations confirm the effectiveness and superiority of our model. However, DEC-Seg encounters challenges in accurately segmenting polyps when dealing with ambiguous areas and unclear boundaries. Some failure cases generated by our model are depicted in Fig. 10. In the $1^{st}$ and $2^{nd}$ rows, while the prominent polyp regions are segmented accurately, there are inaccuracies in parts areas of the segmentation. In the $3^{rd}$ and $4^{th}$ rows, it is evident that the polyp regions are notably small and masked within the background, featuring very unclear boundaries, which presents a significant challenge in accurately segmenting the polyp regions. It is apparent that under these circumstances, our DEC-Seg struggles to identify and segment the polyps. Therefore, dealing with ambiguous areas and segmenting polyps with unclear boundaries is desired to be investigated in future work. Additionally, unlike a simple UNet framework, our model may have higher complexity due to incorporating three decoders and two generative networks. Furthermore, our model has only been validated on 2D segmentation tasks. Moving forward, we intend to address and overcome these limitations in our future research endeavors.
@@ V Conclusion
In this paper, we have presented a novel semi-supervised learning framework (DEC-Seg) for medical image segmentation. The proposed cross-level feature aggregation module integrates the adjacent features from different resolutions, to enhance the features’ representation ability. Then, a scale-enhanced consistency is proposed to handle scale variation and learn more scale-aware features. Meanwhile, we design the scale-fused decoders and a dual-scale complementary fusion module to aggregate the features from the scale-specific decoders and produce the final segmentation maps. Moreover, multiple consistency strategies, i.e., scale-aware perturbation consistency and cross-generative consistency, are presented to enhance the learning process and fully leverage unlabeled data to boost the segmentation performance. Experimental results on multiple datasets from three medical image segmentation tasks show that our DEC-Seg is superior to state-of-the-art semi-supervised segmentation methods.